OpenStudy (anonymous):

Solve each proportion. Leave your answer as a fraction in simplest form. 9/8=k+6/6 (9 over 8 equals K plus 6 over 6)

OpenStudy (anonymous):

\[\frac{9}{8}=\frac{k+6}{6}\iff 6\times 9=8\times (k+6)\]

OpenStudy (anonymous):

What do I do after that part?